OpenStudy (anonymous):

Simplify 8/(2+2i) 2 − 2i 2 + 2i 2 − i 2 + i

myininaya (myininaya):

I suppose that is actually 4/(2-2i) and not 4/2-2i?

OpenStudy (anonymous):

Multiply the denominator and the numerator by the conjugate of the denominator.

OpenStudy (anonymous):

so i would multiply by 2-2i?

OpenStudy (anonymous):

mutiply by the congugate \[\frac{ 8 }{ 2+2i } \times\frac{ 2-2i }{ 2-2i }\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

What celloscope has up there. ^

OpenStudy (anonymous):

From there, it'll be much easier to simplify.

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

ok i got -16-16i/-4-4i^2 now how would i simplify it

OpenStudy (anonymous):

well i^2 simplifies as -1

OpenStudy (anonymous):

and do i divide it by 2

OpenStudy (anonymous):

sryy it took so long to reply.i was eating

OpenStudy (anonymous):

noooo... the demoninator is like this -4-4i^2 so it becomes like this -4-4(-1)

OpenStudy (anonymous):

so -4+4 is on the bottom now.how would i simplify it to one of my answer options?

OpenStudy (anonymous):

@shrutipande9